You wish to invest $600 over one year in two accounts paying 5% and 6% annually. How much should you invest in each to earn $34?
Found 2 solutions by macston, addingup:
Answer by macston(5194)

.
F=amount at 5%; S=amount at 6%
F+S=$600
F=$600-S
0.05F+0.06S=$34 Substitute for F
0.05($600-S)+0.06S=$34
$30-0.05S+0.06S=$34 Subtract $30 from each side.
0.01S=$4 Multiply each side by 100
S=$400 ANSWER 1: You should invest $400 at 6%.
F=$600-S=$600-$400=$200 ANSWER 2: You should invest $200 at 5%.
CHECK:
0.05F+0.06S=$34
0.05($200)+0.06($400)=$34
$10+$24=$34
$34=$34

Answer by addingup(3677)

We are dividing the 600 into two accounts, one pays 5 and the other 6%. So we have:
.06x+.05(600-x)= 34
.06x+30-.05x= 34 On the left subtract x, and subtract 30 on both sides:
.01x= 4 Divide both sides by .01:
x= 400 So you invest 400 at 6% and
600-400= 200 at 5%
Proof:
400*.06= 24
200*.05= 10
Total.....34